What is the final velocity of a 1200 kilogram car starting from rest after 50 meters when 4500 Newton's of force is applied?



1
Expert's answer
2021-05-06T16:52:55-0400

According to the second Newton's law, the acceleration of the car is:

"a=\frac{F}{m}"

F=4500N is total force applied to the car and m=1200kg Is the mass of car

According to the kinematic equation, the distance covered under a constant acceleration is:

"d=v_{0}t+\frac{1}{2}at^2"

v0=0

Since d=50m

"d=\frac{1}{2}at^2"

"t=\sqrt\frac{2d}{a}=\sqrt\frac{2dm}{F}"

The velocity of the car at this moment of time (if it starts from rest) is given by the following kinematic equation:

"v=v_{0}+at"

Substituting the expressions for aa and tt, obtain:

"v=\frac{F}{m}\sqrt\frac{2dm}{F}=\sqrt{\frac{2dF}{m}}"

Put value

"v=\sqrt{\frac{2\times250\times4500}{1200}}=6.1m/sec"

Final velocity of car is 6.1m/sec